1. A cart is pushed up a ramp with a force of 250 N. The ramp sits at a 28° angle with the horizontal.

A. Determine the components of the vector representing the force.

B. Describe what would happen if the angle of the ramp was decreased?



1
Expert's answer
2021-05-24T12:04:56-0400

A. There are two vector components of the force

1.Horizontal component- Horizontal component is the component that acts along

the x axis or we can say along horizontal direction.

Horizontal component "= F\times Cos 28^0=250 \times Cos 28^0=220.74 N"


2.Vertical component-Vertical component is the component that acts along

the y axis or we can say along vertical direction.

Vertical component"=F\times Sin 28^0=250\times Sin28^0 =117.37N"


B. When the angle of the ramp is decreased then the horizontal vector component of the force is increased as the value of "Cosine" increases when the angle is decreased.

But the value of vertical component is decreased as the value of "Sine" decreased when the angle is decreased.